OpenStudy (idealist10):

Solve the initial-value problem (4x^3*y^2-6x^2*y-2x-3)dx+(2x^4*y-2x^3)dy=0, y(1)=3.

OpenStudy (idealist10):

I've got the answer, which is x^4*y^2-2x^3*y-x^2-3x=C but the initial-value is y(1)=3, how do I solve for y?

OpenStudy (idealist10):

@hartnn @myininaya

hartnn (hartnn):

y(1) =3 means when x=1, y=3

OpenStudy (idealist10):

So I found that C=-1 after I plug x=1 and y=3. But how do I solve for y?

hartnn (hartnn):

why do u need y ?

hartnn (hartnn):

its not gonna be easy, if possible

OpenStudy (idealist10):

Because the answer in the book is in the form of y=...

OpenStudy (idealist10):

The answer in the book is...\[y=\frac{ x+\sqrt{2x ^{2}+3x-1} }{ x ^{2} }\]

hartnn (hartnn):

a =x^4 b=-2x^3 c= -x^2-3x+1 then you have ay^2+by+c =0

hartnn (hartnn):

quadratic formula

hartnn (hartnn):

you sure your answer is correct ?

OpenStudy (idealist10):

Well, I guess I'm pretty sure. Let me work the quadratic formula out.

hartnn (hartnn):

i sense you'll get the correct answer :)

OpenStudy (idealist10):

Thank you so much! I got the right answer!